OpenStudy (anonymous):

Evaluate A 2-B 2 for A = 3 and B = -4.

OpenStudy (anonymous):

is that A^2-B^2?

OpenStudy (anonymous):

we have : 3^2-(-4)^2=9-16=-5 Got it ?

OpenStudy (anonymous):

@sione its not -5 its -7 recheck ur math

OpenStudy (anonymous):

Yea ! It s -7 Sorry !

OpenStudy (anonymous):

yep just trying to help in any way